OpenStudy (anonymous):

the cost of 3 markers and 2 pencils is 1.80$ and the cost of 4 markers and 6 pencils is 2.90$ what is the cost of each item?

OpenStudy (anonymous):

solve the eqn

OpenStudy (anonymous):

that doesn't make any sense

OpenStudy (anonymous):

3m+2p=1.8 4m+6p=2.9 -9m-6p=-5.4 -5m=-2.5 marker=$0.5 1.5+2p=1.8 2p=0.3 pencil=$0.15

OpenStudy (anonymous):

Let the number of pencils be x Let the number of markers be y 2x + 3y = 1.80------>1 6x + 4y = 2.90------>2 x = 0.15 y = 0.5 That are the prices, $0.15/pencil and $0.5/marker,